The heat of formation of decane liquid, CioH22, is-60.0 kJ/mol. You place 2.85 g of it into a bomb calorimeter and combust the liquid, forming carbon dioxide gas and liquid water. How many kJ of energy are released (assume a 100% yield)? The calorimeter changes in temperature from 19.50°C to 22.70°C. What is the heat capacity of the calorimeter?
Answer #1

2 C10H22(l) + 31 O2(g) ---> 20 CO2(g) + 22 H2O(l)

DHrxn = (20*DHfCO2 + 22*DHfH2O)-(2*DHfC10H22 + 31*DHfO2)

       = (20*-393.5+ 22*-285.8)-(2*-60+31*0)

       = -14037.6 Kj

DHrxn = -7018.8 kj/mol

no of mol of C10H22 burned = 2.85/142 = 0.02 mol

heat released(q) = 7018.8*0.02 = 140.38 kj

q = c*DT

c = HEAT capacity of calorimeter = ?

DT = 22.7-19.5 = 3.2 c

140.38 = C*3.2

C = 43.87 kj/c